OpenStudy (anonymous):

Multiply. two-thirds times 5 and one-half

OpenStudy (anonymous):

would you like the answer in simplist form ?

OpenStudy (anonymous):

Answer: = 2/3 × 11/2 = (2 × 11) / (3 × 2) = 22/6 = 11/3 = 3 2/3

OpenStudy (anonymous):

3 2/3 is your answer in simplist form

OpenStudy (anonymous):

i hope this helped

OpenStudy (anonymous):

thanks